Solve each equation by the subtraction method. Determine whether the equation are independent, dependant, or inconsistent.
28. x =y + 4
3y – 2y = 4

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE WHICH WILL HELP YOU.IF YOU
STILL HAVE DIFFICULTY PLEASE COME BACK
*******************************************
In the given system of equations is consistent and
dependent, consistent and independent or inconsistent.
3x-y=7
6x-2x=9
what does these terms mean?
1 solutions
--------------------------------------------------------------------------------
Answer 12951 by venugopalramana(684) on 2006-01-13
01:20:47 
IS the given system of equations consistent
MEANS WHETHER THERE IS ANY SET OF VALUES FOR X AND Y
WHICH SATISFIES BOTH THE GIVEN EQUATIONS.
and dependent,MEANS WHETHER ONE EQUATION CAN BE GOT
FROM THE OTHER BY SUITABLE AND PERMISSIBLE
TRANSFORMATION
consistent and independent,MEANS THAT ONE EQUATION
CANNOT BE OBTAINED FROM THE OTHER EQUATION BY BY ANY
SUITABLE AND PERMISSIBLE TRANSFORMATION .
or inconsistent.MEANS THERE ARE NO SET OF VALUES FOR X
AND Y WHICH SATISFY BOTH THE EQUATIONS.
3x-y=7.....I.......
6x-2x=9...I THINK IT SHOULD BE 6X-2Y=9........II
what does these terms mean?NOW I THINK YOU KNOW WHAT
THE TERMS MEAN.THERE ARE TESTS FOR FINDING THESE
SITUATIONS.WE SHALL RESTRICT THE ANALYSIS TO YOUR
PROBLEM AFTER THE CORRECTION MENTIONED ABOVE.
WE FIND THAT ON L.H.S OF THE 2 EQNS.WE HAVE 3X-Y AND
6X-2Y WHICH ARE DIRECT MULTIPLES BY 2 .THAT IS WE GET
LHS OF EQN.2 BY MULTIPLYING THE EQN 1 WITH 2
WE GET EQN.I *2 ....6X-2Y=14
WHERE AS EQN.II SAYS 6X-2Y=9..SO IF WE TRY TO SATISFY
EQN.I ,EQN II WILL NOT BE SATISFIED AND IF WE TRY TO
SATISFY EQN II ,EQN.I WILL NOT BE SATISFIED.SO THERE
ARE NO VALUES OF X AND Y WHICH SATISFY BOTH
EQUATIONS.SO WE SAY THE GIVEN EQNS. ARE INCONSISTENT.